a body of mass 2kg initially at rest on a smooth horizontal plane is subjected to a horizontal forve of magnitude 1/2v+1. find the time when the velocity is 1ms

1
Expert's answer
2022-01-04T16:17:52-0500

"F=ma=m\frac{dv}{dt},"

"\frac 1{2v+1}=2\frac{dv}{dt},"

"(4v+2)dv=dt,"

"t(v)=2v^2+2v,"

"t(1)=4~s."